What volume of a 0.20 m k2SO4 solution contains 57g of k2SO4?

1
Expert's answer
2021-11-29T04:39:25-0500

Using n=CV

V=n/C

n=mass/molar mass

n=57/174

n=0.33

V=0.33/0.2

V=1.65L